When normalizing the parameters for a Cobb Douglas Function, prove that the utility is unique only up to a monotonic transformation if α = 0.3


1
Expert's answer
2021-09-07T19:21:12-0400

Assume the following Cobb Douglas production function

"U= x^\alpha y^{\beta}"

The utility can be unique up to a point where "\alpha+\beta=1"

Since "\alpha" and "\beta" indicate the relative importance of good x and y to a consumer, the Cobb Douglas function can be written as follows.

"U=(X,Y)= x^{\Phi} y^{1-\Phi}"

Where "\Phi= {\alpha \above{2pt} \alpha+\beta}"

and "\Phi= {\beta \above{2pt} \alpha+\beta}"

Now, if "\alpha=0.3" then "\beta=(1-0.3)=0.7"

By nomalizing "\alpha" and "\beta" with "\alpha=0.3" and "\beta=0.7", the utility will be unique upto a monotonic transformation when the Cobb Douglass production function is as follows



"U=x^{0.3} y^{0.7}"